OpenStudy (anonymous):

how do you solve 3500=700*2^8x with logs ?

OpenStudy (shamil98):

3500 = 700(2^8x) 5 = 2^8x log 5 / log 2 = 8x and solve for x from there

OpenStudy (anonymous):

thank you!